In the following circuit, the switch that had been OPEN for a sufficient time was shorted at t=0 In this case, the voltage v(t) at both ends of the capacitor and the current i(t) at the injector are obtained.

1) What is the Capacitor voltage v (0-) value and inductor current i (0-) value at t=0- just before t=0-

2) If sufficient time has passed since switching occurred, what is the value of v(00) and i(00) when t=oo?

3) In the case of t t0 in the RLC circuit above, obtain a characteristic equation for v(t) and obtain the two years. Use the results to determine which of these RLC circuits is under-, critical- or over-damping.

4) If t t0, obtain a v(t) expression.

switch to open 42 THI eu w 32n2 TSEV @ 12V tot
Answer #1

In the case of at t= to in the RLC citant above obtained a characteristic eaucation Hove vet) and obtain the two yeats. UseVes a Schlady state xesponse tu - 1.0206 a 21 S2 + 55+ 6zo 92+24 wnist wie Wn = ato 5 2 Cu xao = 5, 206 rá tu= 1 critically d42 АН t je + 12V 22. •5 F E det IVCA = -2 À ë 20 - 3B e 30 c d = je du je Apply kel at node- ů (ot) = je (ot) + Xcot) o Jelot